A chain with negligible mass is wrapped around a g
Answer #1

here,

R = 2 m

m = 14 kg

height descends , h = 10 m

t = 2 s

let the accelration be a

h = u*t + 0.5 * a * t^2

10 = 0 + 0.5 * a * 2^2

a = 5 m/s^2

let the mass of the gear be M

a = ( net force) /( effective mass)

a = ( m * g) /( m + (I) /R^2)

5 = (14 * 9.8 ) /( 14 + 0.5 * M)

M = 26.88 kg

the moment of inertia of gear , I = 0.5 * M * R^2

I = 0.5 * 26.88 * 2^2

I = 53.76 kg.m/s^2

the moment of inertia of the gear is 53.76 kg.m/s^2